OpenStudy (anonymous):

simplify: (1/4 a^2b^-3)^-1/4

OpenStudy (gorv):

\[(\frac{ 1 }{ 4 }*a^{2}*b^{-3})^{-1/4}\]

OpenStudy (mathmale):

If possible, please use Equation Editor for greater clarity. @gorv's representation of your algebraic expression is much easier to read.

OpenStudy (gorv):

@mathmale but it is showing very small size

OpenStudy (anonymous):

that is the correct problem. where do I find equation editor

OpenStudy (mathmale):

Look for the blue button at the very bottom of your OpenStudy screen:\[\sum_{}^{ }Equation\]

OpenStudy (gorv):

in bottom of comment box at left corner its written

OpenStudy (anonymous):

(1/4 a^2b^-3)^-1/4=\[(\sqrt[4]{(1/4 a^2b^-3)})^-1=1/\sqrt[4]{(1/4 a^2b^-3)}\] =1 / 1/2 ^2/4 a * b^-3/4 = 1/1/2^-2a*b-3/4=1/4a * \[\sqrt[4]{1/b^3}\] I think !:)

OpenStudy (anonymous):

thanks but that's not one of my choices

OpenStudy (anonymous):

What are the choices ?

OpenStudy (anonymous):

sq rt. 2a^1/2b3/4 (sq rt 2/a^1/2 b^3/4) (a^1/2)/sq rt 2b ^3/4 s(sq rt 2b ^3/4)/a^1/2

OpenStudy (anonymous):

Is this your question ?: \[((1/4 a^2) (* b^-3))^-1/4\]

OpenStudy (anonymous):

@berrypicking

OpenStudy (anonymous):

yes